# On the minimisation of the Peak–to–average ratio

### v1

@@ Abstract.
$\ \ \ \ \ \mathrm{E}_{p}(u):=\big{\|}\mathrm{H}\big{(}\cdot,u,\mathrm{D}u,\ldots,\mathrm{D}^{k}u\big{)}\big{\|}_{\mathrm{L}^{p}(\Omega)},\ \ \ u\in\mathrm{W}^{k,p}(\Omega;\mathbb{R}^{N}).$
@@
$\ \ \ \ \ \mathrm{C}_{\infty,p}(u_{\infty})\,=\,\inf\Big{\{}\mathrm{C}_{\infty,p}(u)\ :\ u\in\mathrm{W}^{k,\infty}_{\varphi}(\Omega;\mathbb{R}^{N}),\ \mathrm{E}_{1}(u)\neq 0\Big{\}},$
@@
$\left\{\ \ \begin{array}[]{ll}\big{|}\mathrm{H}\big{(}\cdot,u,\mathrm{D}u,\ldots,\mathrm{D}^{k}u\big{)}\big{|}=\Lambda,&\text{ a.e.\ in }\Omega,\\ u=\varphi,&\text{ on }\partial\Omega,\\ \mathrm{D}u=\mathrm{D}\varphi,&\text{ on }\partial\Omega,\\ \ \ \ \ \vdots&\ \ \ \ \ \vdots\\ \mathrm{D}^{k-1}u=\mathrm{D}^{k-1}\varphi,&\text{ on }\partial\Omega.\end{array}\right.$
@@ 1. Introduction
(1.0.1) $\ \ \ \ \ \mathrm{E}_{p}(u):=\big{\|}\mathrm{H}(\mathrm{D}^{[k]}u)\big{\|}_{\mathrm{L}^{p}(\Omega)},\ \ \ u\in\mathrm{W}^{k,p}(\Omega;\mathbb{R}^{N}).$
@@
$\|f\|_{\mathrm{L}^{p}(\Omega)}:=\left(\,-\displaystyle\int_{\Omega}|f|^{p}\,\mathrm{d}\mathcal{L}^{n}\!\right)^{\!1/p}.$
@@
(1.0.2) $\ \ \ \ \ \ \ \mathrm{C}_{\infty,p}(u)\,:=\,\frac{\mathrm{E}_{\infty}(u)}{\mathrm{E}_{p}(u)},\ \ \ \ u\in\mathrm{W}^{k,\infty}(\Omega;\mathbb{R}^{N}),\ \mathrm{E}_{1}(u)\neq 0.$
@@
(1.0.3) $\ \ \ \ \ \mathrm{C}_{\infty,p}(u_{\infty})\,=\,\inf\Big{\{}\mathrm{C}_{\infty,p}(u)\ :\ u\in\mathrm{W}^{k,\infty}_{\varphi}(\Omega;\mathbb{R}^{N}),\ \mathrm{E}_{1}(u)\neq 0\Big{\}},$
@@
In other words, we are interested in finding a minimiser $u_{\infty}\in\mathrm{W}^{k,\infty}_{\varphi}(\Omega;\mathbb{R}^{N})$ which realises the infimum in (1.0.3). The primary motivation to study the problem (1.0.3) comes from applications. When $p\in\{1,2\}$ in (1.0.2), the functionals $\mathrm{C}_{\infty,1}$ and $\mathrm{C}_{\infty,2}$ are known as the “Crest factor” and arises as the “peak–to–average ratio” problem in numerous applications, including eg. the modelling of nuclear reactors, engineering, signal and sound processing, se e.g. [15, 16, 17, 24, 26, 28]. Further, if $\mathrm{H}(x,\cdot,\cdot)$ is for all $x\in\Omega$ positively homogeneous of some degree $K>0$, then (1.0.3) is equivalent the the following constrained $\mathrm{L}^{\infty}$ variational problem (with integral constraint):
(1.0.4) $\ \ \ \ \ \mathrm{E}_{\infty}(u_{\infty})\,=\,\inf\Big{\{}\mathrm{E}_{\infty}(u)\ :\ u\in\mathrm{W}^{k,\infty}_{\varphi}(\Omega;\mathbb{R}^{N}),\ \mathrm{E}_{p}(u)=1\Big{\}}.$
@@
(1.0.5) $\left\{\ \ \ \begin{array}[]{ll}\big{|}\mathrm{H}\big{(}\cdot,u,\mathrm{D}u,\ldots,\mathrm{D}^{k}u\big{)}\big{|}=\Lambda,&\text{ a.e.\ in }\Omega,\\ u=\varphi,&\text{ on }\partial\Omega,\\ \mathrm{D}u=\mathrm{D}\varphi,&\text{ on }\partial\Omega,\\ \ \ \ \ \vdots&\ \ \ \ \vdots\\ \mathrm{D}^{k-1}u=\mathrm{D}^{k-1}\varphi,&\text{ on }\partial\Omega.\end{array}\right.$
Given our notation for jets, we can rewrite the $(k-1)$-th order Dirichlet problem (1.0.5) in the following abbreviated form as:
@@ 2. Main results and proofs
    $f(\textbf{X})\,\leq\,-\displaystyle\int_{[0,1]^{n}}f\big{(}\textbf{X}+\mathrm{D}^{k}\psi\big{)}\,\mathrm{d}\mathcal{L}^{n},$
[…]
@@
(2.0.4) $\left\{\ \ \ \begin{split}&\text{For any $x\in\Omega$ and $X$ in a bounded subset of $\mathbb{R}^{N}\times\mathbb{R}^{Nn}\times\cdots\times\mathbb{R}^{Nn^{(k-1)}}_{s}$,}\\ &\ \text{ there exists }\Lambda\geq\Lambda_{*},\text{ such that }\big{\{}|\mathrm{H}(x,X,\cdot)|\leq\Lambda\}\text{ is bounded in }\mathbb{R}^{Nn^{k}}_{s}.\end{split}\right.$
@@
(2.0.5) $\left\{\ \ \ \begin{split}&\text{For any $x\in\Omega$, $X$ in a bounded subsetof $\mathbb{R}^{N}\times\mathbb{R}^{Nn}\times\cdots\times\mathbb{R}^{Nn^{(k-1)}}_{s}$}\\ &\ \,\text{ and any }\Lambda\geq\Lambda_{*},\text{ the set }\big{\{}|\mathrm{H}(x,X,\cdot)|\leq\Lambda\}\text{ is bounded in }\mathbb{R}^{Nn^{k}}_{s}.\end{split}\right.$
@@
(2.0.8) $\begin{cases}\ \ \mathrm{H}(\cdot,u,\mathrm{D}u)=\Lambda,\ &\text{a.e.\ on }\Omega,\\ \ \ u=\varphi,\ &\text{on }\partial\Omega,\\ \ \ \mathrm{D}u=\mathrm{D}\varphi,\ &\text{on }\partial\Omega,\end{cases}$
@@
Finally, the result above remains true if $h\geq 0$, but we only require that $t\mapsto h\big{(}x,X,t\hskip 1.0pt\mathbb{I}_{n}\big{)}$ be strictly increasing on $[0,\infty)$ (alongside (2.0.6) and (2.0.7)), resulting in the exact same conclusions via the same proof, with the additional premise that $\Lambda_{*}\geq 0$.
@@
For any matrix $\mathrm{X}\in\mathbb{R}^{n^{2}}_{s}$, let $\{\lambda_{1}(\mathrm{X})$, … ,$\lambda_{n}(\mathrm{X})\}$ denote its eigenvalues in increasing order. Consider the following Dirichlet problem involving singular values of the gradient matrix
(2.0.14) $\ \ \ \begin{cases}\ \ \lambda_{i}(\mathrm{D}u^{\top}\mathrm{D}u)\,=\,\alpha(\cdot,u),&\text{a.e. on }\Omega,\quad i=1,\ldots,n,\\ \ \ u=\varphi,\,&\text{on }\partial\Omega,\\ \ \ \mathrm{D}u=\mathrm{D}\varphi,&\text{on }\partial\Omega.\end{cases}$
@@
a given continuous function. Let $\mathrm{C}_{\infty,p}$ be the Crest factor functional given by (1.0.1)-(1.0.2), and fix $\varphi\in\mathrm{W}^{k,\infty}(\Omega;\mathbb{R}^{N})$. Suppose further that the problem (1.0.5) is solvable [^1] Propositions 1 and 2 provide sufficient conditions on $\mathrm{H},\varphi,\Lambda_{*}$ for this to be true. for all $\Lambda\geq\Lambda_{*}$, for some $\Lambda_{*}\geq 0$.
@@
    There exists $\Lambda\geq\Lambda_{*}$, such that $u_{\infty}$ is a strong a.e. solution to the Dirichlet problem (1.0.5), namely
@@
Proof of Theorem 3. (2) $\Rightarrow$ (1): Suppose that $(u_{\infty},\Lambda)\in\mathrm{W}^{k,\infty}_{\varphi}(\Omega;\mathbb{R}^{N})\times[\Lambda_{*},\infty)$ solves the problem (1.0.5), where $\Lambda_{*}\geq 0$ is given. This implies that $\big{|}\mathrm{H}\big{(}\mathrm{D}^{[k]}u_{\infty}\big{)}\big{|}=\Lambda$ a.e. on $\Omega$, and also that $u_{\infty}-\varphi\in\mathrm{W}^{k,\infty}_{0}(\Omega;\mathbb{R}^{N})$, or equivalently
@@
Note that, since we are using the rescaled $L^{p}$ norms on $\Omega$, Hölder’s inequality implies that
@@
which implies that the $(k-1)$-th order Dirichlet boundary condition of (1.0.5) is satisfied. Further, $\mathrm{E}_{1}(u_{\infty})\neq 0$. Equivalently, $\mathrm{H}(\mathrm{D}^{[k]}u_{\infty})\neq 0$ on a Borel subset of positive Lebesgue measure on $\Omega$. By Hölder’s inequality, we again have
@@
(2.0.17) $\mathrm{C}_{\infty,p}(u)\,\geq\,1,\ \text{ for all }u\in\mathrm{W}^{k,\infty}_{\varphi}(\Omega;\mathbb{R}^{N}),\ \text{when }\mathrm{E}_{1}(u_{\infty})\neq 0.$
@@
To this aim, fix any solution $(u_{0},\Lambda)\in\mathrm{W}^{k,\infty}_{\varphi}(\Omega;\mathbb{R}^{N})\times[\Lambda_{*},\infty)$ to (1.0.5), guaranteed to exist for some $\Lambda_{*}\geq 0$ by our assumptions. The arguments used above in the implication “(2) $\Rightarrow$ (1)” (with $u_{0}$ in the place of $u_{\infty}$) prove that $u_{0}$ is a global minimiser with energy value equal to $1$, namely
@@
(2.0.18) $\mathrm{E}_{\infty}(u_{\infty})^{p}\,=\,-\displaystyle\int_{\Omega}\big{|}\mathrm{H}\big{(}\mathrm{D}^{[k]}u_{\infty}\big{)}\big{|}^{p}\,\mathrm{d}\mathcal{L}^{n}.$
@@
(2.0.19) $\Omega\,=\,\Big{\{}\big{|}\mathrm{H}\big{(}\mathrm{D}^{[k]}u_{\infty}\big{)}\big{|}\,=\,\mathrm{E}_{\infty}(u_{\infty})\Big{\}}\bigcup\Big{\{}\big{|}\mathrm{H}\big{(}\mathrm{D}^{[k]}u_{\infty}\big{)}\big{|}\,<\,\mathrm{E}_{\infty}(u_{\infty})\Big{\}}.$
@@
$\begin{split}\mathrm{E}_{\infty}(u_{\infty})^{p}\,=&\ -\displaystyle\int_{\Omega}\big{|}\mathrm{H}\big{(}\mathrm{D}^{[k]}u_{\infty}\big{)}\big{|}^{p}\,\mathrm{d}\mathcal{L}^{n}\\ =&\ \frac{1}{\mathcal{L}^{n}(\Omega)}\int_{\big{\{}|\mathrm{H}(\mathrm{D}^{[k]}u_{\infty})|=\mathrm{E}_{\infty}(u_{\infty})\big{\}}}\big{|}\mathrm{H}\big{(}\mathrm{D}^{[k]}u_{\infty}\big{)}\big{|}^{p}\,\mathrm{d}\mathcal{L}^{n}\\ &\ +\frac{1}{\mathcal{L}^{n}(\Omega)}\int_{\big{\{}|\mathrm{H}(\mathrm{D}^{[k]}u_{\infty})|<\mathrm{E}_{\infty}(u_{\infty})\big{\}}}\big{|}\mathrm{H}\big{(}\mathrm{D}^{[k]}u_{\infty}\big{)}\big{|}^{p}\,\mathrm{d}\mathcal{L}^{n}\\ =&\ \frac{\mathrm{E}_{\infty}(u_{\infty})^{p}}{\mathcal{L}^{n}(\Omega)}\mathcal{L}^{n}\Big{(}\Big{\{}|\mathrm{H}(\mathrm{D}^{[k]}u_{\infty})|=\mathrm{E}_{\infty}(u_{\infty})\Big{\}}\Big{)}\\ &\ +\frac{1}{\mathcal{L}^{n}(\Omega)}\int_{\big{\{}|\mathrm{H}(\mathrm{D}^{[k]}u_{\infty})|<\mathrm{E}_{\infty}(u_{\infty})\big{\}}}\big{|}\mathrm{H}\big{(}\mathrm{D}^{[k]}u_{\infty}\big{)}\big{|}^{p}\,\mathrm{d}\mathcal{L}^{n}.\end{split}$
@@
$\begin{split}-\displaystyle\int_{\big{\{}|\mathrm{H}(\mathrm{D}^{[k]}u_{\infty})|<\mathrm{E}_{\infty}(u_{\infty})\big{\}}}\big{|}\mathrm{H}\big{(}\mathrm{D}^{[k]}u_{\infty}\big{)}\big{|}^{p}\,\mathrm{d}\mathcal{L}^{n}\ =\,\mathrm{E}_{\infty}(u_{\infty})^{p}.\end{split}$
@@
$\Lambda:=\mathrm{E}_{\infty}(u_{\infty})\geq 0.$

### v2

@@ Abstract.
[…]
@@
[…]
@@
[…]
@@ 1. Introduction
[…]
@@
$\|f\|_{\mathrm{L}^{p}(\Omega)}:=\left(\,-\hskip-13.0pt\displaystyle\int_{\Omega}|f|^{p}\,\mathrm{d}\mathcal{L}^{n}\!\right)^{\!1/p}.$
@@
[…]
@@
[…]
@@
In other words, we are interested in finding a minimiser $u_{\infty}\in\mathrm{W}^{k,\infty}_{\varphi}(\Omega;\mathbb{R}^{N})$ which realises the infimum in (1.0.3). The primary motivation to study the problem (1.0.3) comes from applications. When $p\in\{1,2\}$ in (1.0.2), the functionals $\mathrm{C}_{\infty,1}$ and $\mathrm{C}_{\infty,2}$ are known as the “Crest factor” and arises as the “peak–to–average ratio” problem in numerous applications, including eg. the modelling of nuclear reactors, engineering, signal and sound processing, se e.g. [15, 16, 17, 24, 26, 28]. Further, if $\mathrm{H}(x,\cdot,\cdot)$ is for all $x\in\Omega$ positively homogeneous of some degree $K>0$, then (1.0.3) is equivalent to the following constrained $\mathrm{L}^{\infty}$ variational problem (with integral constraint):
[…]
@@
[…]
Given our notation for jets, we can rewrite the (k1)(k-1)-th order Dirichlet problem (1.0.5) in the following abbreviated form as:
@@ 2. Main results and proofs
    $f(\textbf{X})\,\leq\,-\hskip-13.0pt\displaystyle\int_{[0,1]^{n}}f\big(\textbf{X}+\mathrm{D}^{k}\psi\big)\,\mathrm{d}\mathcal{L}^{n},$
    $\text{ for all }\psi\in\mathrm{W}^{k,\infty}_{0}({\Omega};\mathbb{R}^{N})\text{ and all }\textbf{X}\in\mathbb{R}^{Nn^{k}}_{s}.$
@@
[…]
@@
[…]
@@
(2.0.8) $\begin{cases}\ \ \mathrm{H}(\cdot,u,\mathrm{D}u)=\Lambda,\ &\text{a.e.\ on }\Omega,\\ \ \ u=\varphi,\ &\text{on }\partial\Omega,\end{cases}$}
@@
Finally, the result above remains true if $h\geq 0$, but we only require that $t\mapsto h\big(x,X,t\hskip 1.0pt\mathbb{I}_{n}\big)$ be strictly increasing on $0,\infty)$ (alongside (2.0.6) and (2.0.7)), resulting in the exact same conclusions via the same proof, with the additional premise that $\Lambda_{*}\geq 0$.
@@
For any matrix $\mathrm{X}\in\mathbb{R}^{n^{2}}_{s}$, let $\{\lambda_{1}(\mathrm{X})$, … $\lambda_{n}(\mathrm{X})\}$ denote its eigenvalues in increasing order. Consider the following Dirichlet problem involving singular values of the gradient matrix
[…]
@@
a given continuous function. Let $\mathrm{C}_{\infty,p}$ be the Crest factor functional given by (1.0.1)-(1.0.2), and fix $\varphi\in\mathrm{W}^{k,\infty}(\Omega;\mathbb{R}^{N})$. Suppose further that the problem (1.0.5) is solvable [^1] Propositions 1 and 2 provide sufficient conditions on $\mathrm{H},\varphi,\Lambda_{*}$ for this to be true. for all $\Lambda\geq\Lambda_{*}$, for some $\Lambda_{*}>0$.
@@
    There exists $\Lambda>0$, such that $u_{\infty}$ is a strong a.e. solution to the Dirichlet problem (1.0.5), namely
@@
Proof of Theorem 3. (2) $\Rightarrow$ (1): By assumption, we have that $(u_{\infty},\Lambda)\in\mathrm{W}^{k,\infty}_{\varphi}(\Omega;\mathbb{R}^{N})\times(0,\infty)$ solves the problem (1.0.5). This implies that $\big|\mathrm{H}\big(\mathrm{D}^{[k]}u_{\infty}\big)\big|=\Lambda$ a.e. on $\Omega$, and also that $u_{\infty}-\varphi\in\mathrm{W}^{k,\infty}_{0}(\Omega;\mathbb{R}^{N})$, or equivalently
@@
In particular, the satisfaction of the PDE implies
$\Lambda=\underset{\Omega}{\mathrm{ess}\sup}\,\big|\mathrm{H}\big(\mathrm{D}^{[k]}u_{\infty}\big)\big|=\mathrm{E}_{\infty}(u_{\infty}),$
and also
$\Lambda=\left(\,-\hskip-13.0pt\displaystyle\int_{\Omega}\big|\mathrm{H}\big(\mathrm{D}^{[k]}u_{\infty}\big)\big|^{p}\,\mathrm{d}\mathcal{L}^{n}\!\right)^{\!1/p}=\mathrm{E}_{p}(u_{\infty}),$
for any $p\geq 1$. Note that, since we are using the rescaled $L^{p}$ norms on $\Omega$, Hölder’s inequality implies that
@@
which implies that the (k1)(k-1)-th order Dirichlet boundary condition of (1.0.5) is satisfied. Further, $\mathrm{E}_{1}(u_{\infty})\neq 0$. Equivalently, $\mathrm{H}(\mathrm{D}^{[k]}u_{\infty})\neq 0$ on a Borel subset of positive Lebesgue measure on $\Omega$, which in turn gives that $\mathrm{E}_{\infty}(u_{\infty})>0$. By Hölder’s inequality, we again have
@@
To this aim, fix any solution $(u_{0},\Lambda_{0})\in\mathrm{W}^{k,\infty}_{\varphi}(\Omega;\mathbb{R}^{N})\times[\Lambda_{*},\infty)$ to (1.0.5), guaranteed to exist for some $\Lambda_{*}>0$ by our assumptions. The arguments used above in the implication “(2) $\Rightarrow$ (1)” (with $u_{0}$ in the place of $u_{\infty}$) prove that $u_{0}$ is a global minimiser with energy value equal to $1$, namely
@@
(2.0.18) $\mathrm{E}_{\infty}(u_{\infty})^{p}\,=\,-\hskip-13.0pt\displaystyle\int_{\Omega}\big|\mathrm{H}\big(\mathrm{D}^{[k]}u_{\infty}\big)\big|^{p}\,\mathrm{d}\mathcal{L}^{n}.$
@@
(2.0.19) $\mathcal{L}^{n}\bigg(\Omega\setminus\Big[\Big\{\big|\mathrm{H}\big(\mathrm{D}^{[k]}u_{\infty}\big)\big|\,=\,\mathrm{E}_{\infty}(u_{\infty})\Big\}\bigcup\Big\{\big|\mathrm{H}\big(\mathrm{D}^{[k]}u_{\infty}\big)\big|\,<\,\mathrm{E}_{\infty}(u_{\infty})\Big\}\Big]\bigg)=0.$}
@@
$\begin{split}\mathrm{E}_{\infty}(u_{\infty})^{p}\,=&\ -\hskip-13.0pt\displaystyle\int_{\Omega}\big|\mathrm{H}\big(\mathrm{D}^{[k]}u_{\infty}\big)\big|^{p}\,\mathrm{d}\mathcal{L}^{n}\\ =&\ \frac{1}{\mathcal{L}^{n}(\Omega)}\int_{\big\{|\mathrm{H}(\mathrm{D}^{[k]}u_{\infty})|=\mathrm{E}_{\infty}(u_{\infty})\big\}}\big|\mathrm{H}\big(\mathrm{D}^{[k]}u_{\infty}\big)\big|^{p}\,\mathrm{d}\mathcal{L}^{n}\\ &\ +\frac{1}{\mathcal{L}^{n}(\Omega)}\int_{\big\{|\mathrm{H}(\mathrm{D}^{[k]}u_{\infty})|<\mathrm{E}_{\infty}(u_{\infty})\big\}}\big|\mathrm{H}\big(\mathrm{D}^{[k]}u_{\infty}\big)\big|^{p}\,\mathrm{d}\mathcal{L}^{n}\\ =&\ \frac{\mathrm{E}_{\infty}(u_{\infty})^{p}}{\mathcal{L}^{n}(\Omega)}\mathcal{L}^{n}\Big(\Big\{|\mathrm{H}(\mathrm{D}^{[k]}u_{\infty})|=\mathrm{E}_{\infty}(u_{\infty})\Big\}\Big)\\ &\ +\frac{1}{\mathcal{L}^{n}(\Omega)}\int_{\big\{|\mathrm{H}(\mathrm{D}^{[k]}u_{\infty})|<\mathrm{E}_{\infty}(u_{\infty})\big\}}\big|\mathrm{H}\big(\mathrm{D}^{[k]}u_{\infty}\big)\big|^{p}\,\mathrm{d}\mathcal{L}^{n}.\end{split}$
@@
$\begin{split}-\hskip-13.0pt\displaystyle\int_{\big\{|\mathrm{H}(\mathrm{D}^{[k]}u_{\infty})|<\mathrm{E}_{\infty}(u_{\infty})\big\}}\big|\mathrm{H}\big(\mathrm{D}^{[k]}u_{\infty}\big)\big|^{p}\,\mathrm{d}\mathcal{L}^{n}\ =\,\mathrm{E}_{\infty}(u_{\infty})^{p}.\end{split}$
@@
$\Lambda:=\mathrm{E}_{\infty}(u_{\infty})>0$}.
@@
Remark 4.
In the context of Theorem 3 above, it might be reasonable to expect that $\Lambda\geq\Lambda_{*}$ in part (2), for the same $\Lambda_{*}$ as in the assumptions. This seems generically speaking to be the case, but it does not appear directly deducible without imposing specific assumptions on $\mathrm{H}$ (instead of relying on the weaker requirement of mere solvability of (1.0.5)). Let us demonstrate this in a simple particular case. If for example $|\mathrm{H}|$ is $k$-quasiconvex depending only on the leading term of $k$-th order derivatives, and the boundary condition $\phi$ is a polynomial of degree $k$, then as in Proposition 1 one can define $\Lambda_{*}$ by (2.0.2). In that case, by Hölder’s inequality, the definition of $k$-quasiconvexity, the Gauss-Green theorem and the satisfaction of the Dirichlet boundary condition up to derivatives of order k1k-1, we can indeed estimate
$\begin{split}\Lambda=\underset{\Omega}{\mathrm{ess}\,\sup}\,\big|\mathrm{H}(\mathrm{D}^{k}u_{\infty})\big|&\geq-\hskip-13.0pt\displaystyle\int_{\Omega}\big|\mathrm{H}(\mathrm{D}^{k}u_{\infty})\big|\,\mathrm{d}\mathcal{L}^{n}\\ &\geq\left|\mathrm{H}\!\left(\,-\hskip-13.0pt\displaystyle\int_{\Omega}\mathrm{D}^{k}u_{\infty}\,\mathrm{d}\mathcal{L}^{n}\right)\right|\\ &=\left|\mathrm{H}\!\left(\,-\hskip-13.0pt\displaystyle\int_{\Omega}\mathrm{D}^{k}\phi\,\mathrm{d}\mathcal{L}^{n}\right)\right|\\ &=\underset{\Omega}{\mathrm{ess}\,\sup}\,\big|\mathrm{H}(\mathrm{D}^{k}\phi)\big|\\ &\geq\Lambda_{*}.\end{split}$